Maya and Rena bought a 60 minute phone card to call their mom in Ghana. Maya used three times as many minutes as Rena. How many minutes did each girl use?
Answer by JulietG(1812)

M + R = 60
M = 3R
... (3R) + R = 60
4R = 60
R = 15
Rena used 15 minutes. Therefore Maya used 45 minutes (3 times as many)
15+45=60